OpenStudy (anonymous):

the circle x^2 + y^2 = 49 consists of four one-to-one functions f1, f2, f3, f4. (1) find a formula for the function f1 and its domain. (2) find the inverse of function f1 and write it as y= f^(-1)x

OpenStudy (anonymous):

you could say \[f_1(x)=\sqrt{1-x^2}\text{ for } 0\leq x<1\] for example

OpenStudy (anonymous):

why is that?

OpenStudy (anonymous):

you want a one to one function so you have to break it into 4 parts 

OpenStudy (anonymous):

there are 4 one to one functions that together give the circle.

OpenStudy (anonymous):

if you label them \[f_1,f_2,f_3,f_4\] respectively, they will give you the whole thing. you cut them off by restricting the domain

OpenStudy (anonymous):

you can see that \[f_1\] is for x from 0 to 1

OpenStudy (anonymous):

wouldnt it be f(x) = x + 7 because you just solve for y and it's positive in the first quadrant?

OpenStudy (anonymous):

the answer you gave me is apparently wrong

OpenStudy (anonymous):

ok let me do it correctly. the radius is 7, and so it would be \[f_1(x)=\sqrt{49-x^2} \text{ for } 0\leq x < 7\]

OpenStudy (anonymous):

i forgot the radius was 7 and not 1, sorry

OpenStudy (anonymous):

thats alright and thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

to find the inverse do i just solve for x from that equation?